What is the hybridisation of Phosphorus in PH3?

Grade:8

2 Answers

Sunil Kumar FP
9 years ago
for phospine it is sp3 because the hybrid was just resulted from sigma bond of the 3s orbital of the H and p orbital of the P,however, only 3 bonds will form because the remaining two electrons will be a lone pair
for hydrogen sulfide
hybridisation=3+1/2(5-3)=4=sp^3
=no of sorrounding atom +1/2(group number-valency of sorrounding atom)
vineet chatterjee
9 years ago
sir,will the drago rule not be applied here?